# Running in circles: practical limitations for real-life application of data fission and data thinning in post-clustering differential analysis

### v2

@@ 2.1 Data fission and data thinning
2.2 Limits in practical application of data fission and data thinning
Data fission and data thinning methods face a circular challenge when applied for post-clustering inference to mixture distributions used to describe clustered data. They require accurate estimates of intra-component parameters (like the variance), which depend on knowing the true cluster – but estimating the clusters is the whole point of the analysis in the first place, and true clusters are never known a priori.
2.2.1 Mixture distributions
@@
Figure 1: Schematic view illustrating the circularity induced by data fission for post-clustering differential analysis
Data fission and data thinning therefore create a circular situation, as illustrated in the schematic Figure 1. To be applied for post-clustering inference, these methods require knowledge of the intra-component parameters $\boldsymbol{\theta}_{k}$ which, in turn, depend themselves on the components that we would estimate. An effective approach could be to estimate a global parameter $\hat{\boldsymbol{\theta}}$ based on all observations. However, this would make the critical assumption that the parameter value is the same across all components, i.e. $\boldsymbol{\theta}=\boldsymbol{\theta}_{k}$; this means that the data are distributed according to the same distribution globally, regardless of their component:
$p(\mathbf{x}_{i})=f(\mathbf{x}_{i}|\boldsymbol{\theta}_{k})=f(\mathbf{x}_{i}|\boldsymbol{\theta}).$ (1)
This corresponds to the global null of no cluster. Yun and Foygel Barber, (2023) highlighted similar challenges for existing post-clustering selective tests.
2.2.2 Scale parameter prior knowledge and estimation
We will restrict our analysis to the Gaussian setting, that is considering:
$f\left(\mathbf{x}_{i}|\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\right)=\frac{1}{(2\pi)^{p/2}|\boldsymbol{\Sigma}_{g}|^{1/2}}\exp\left(-\frac{1}{2}(\mathbf{x}_{i}-\boldsymbol{\mu}_{k})^{T}\boldsymbol{\Sigma}_{k}^{-1}(\mathbf{x}_{i}-\boldsymbol{\mu}_{k})\right)$
[…]
Figure 2: Toy example illustrating the impact of variance estimation on data fission p-values Panel A: A two-dimensional Gaussian distribution incorrectly clustered into $2$ clusters. Panel B: QQ-plot of the $t$-test p-values for the comparison between the two estimated clusters across $1,\!000$ simulations when data fission is performed with $3$ different variance estimators. Panel C: Extension of the problem to a two-component, two-dimensional Gaussian mixture incorrectly clustered into $3$ clusters. $C_{1}$ and $C_{2}$ originate from the same component, which is erroneously split into two. Panel D: $t$-test p-values for the comparison between $C_{1}$ and $C_{2}$ over $1,\!000$ data simulations using the same three variance estimators for the data fission.
Figure 2 provides a comprehensive overview of the challenges associated with variance estimation in data fission. Panel A presents an illustrative example involving $300$ realizations of a multivariate Gaussian distribution ($K=1$) with a mean vector $\boldsymbol{\mu}=\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right)$ and a covariance matrix $\boldsymbol{\Sigma}=\left(\begin{smallmatrix}4&0\\ 0&1\end{smallmatrix}\right)$. The $k-$means algorithm was applied to this dataset, resulting in the estimation of two clusters, $C_{1}$ and $C_{2}$. Data fission performance was evaluated when using different estimations of $\boldsymbol{\Sigma}$. First, we considered the true intra-components covariance matrix $\boldsymbol{\Sigma}_{k}$ (see section 2.3.1 for a proposed approach to data fission assuming several mixture component). We also considered the overall sample covariance matrix $\widehat{\boldsymbol{\Sigma}}=\frac{1}{n-1}\sum_{i=1}^{n}\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}\right)\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}\right)^{t}$ where $\overline{\boldsymbol{X}}$ is the sample mean vector. We finally use the $k-$means results to compute an intra-cluster covariance matrix defined as: $\widehat{\boldsymbol{\Sigma}}_{\hat{k}}=\frac{1}{|C_{k}|-1}\sum_{i=1\in C_{k}}\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}_{C_{k}}\right)\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}_{C_{k}}\right)^{t}$ where $\overline{\boldsymbol{X}}_{C_{k}}$ is the sample mean vector of the cluster $C_{k}$. Since both clusters originate from the same component, there is no inherent differences between them, implying that the $t$-test p-values should exhibit a uniform distribution. Panel B presents a QQ-plot illustrating the resulting p-values against the Uniform distribution for the test on $X_{1}$ in $1,\!000$ replications of the experiment. In this scenario where the mixture has only one component ($K=1$), the sample covariance matrix $\widehat{\boldsymbol{\Sigma}}$ provides an unbiased estimate of the true $\boldsymbol{\Sigma}$, resulting in uniformly distributed p-values. However, when considering the intra-cluster covariance matrices $\widehat{\boldsymbol{\Sigma}}_{\hat{k}}$ (for $k=1,2$) derived from the $k-$means results, the p-values no longer exhibit a uniform distribution. In this case, the estimated matrices $\widehat{\boldsymbol{\Sigma}}_{\hat{k}}$ for each cluster drastically underestimate the true covariance matrix $\boldsymbol{\Sigma}$, compromising the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$. This deviation from independence leads to spuriously concluding the estimated clusters are truly different from one another.
Panel C introduces a scenario with two true clusters generated using a mixture of two Gaussian distributions ($K=2$). The $k-$means identifies $3$ clusters (for illustration purposes), incorrectly splitting one mixture component into clusters $C_{1}$ and $C_{2}$. Data fission performance was again compared when the decomposition is performed using the same $3$ covariance estimators as in the previous scenario. Panel D presents the resulting p-values for the $t$-test on $C_{1}$ and $C_{2}$ on $X_{1}$ over $1,\!000$ replications of the analysis. Efficient Type I error control is achievable only by considering the true intra-components covariance matrix. Both global and estimated intra-cluster covariances are biased estimators, leading to correlations between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, compromising control of the Type I error rate. This underscores the intricate challenges associated with covariance estimation for data fission in practical scenarios, mainly due to the misspecification of the generative model for data fission arising when genuine clusters exist within the dataset.
2.3 Practical solutions
To address the circular situation of data fission and thinning in Gaussian mixture models, we propose modeling each observation as a realization of its own Gaussian distribution. This approach bypasses the need for prior knowledge of the true data structure, allowing for individual-level data fission and thinning. Nonetheless, the accurate estimation of individual variance parameters remains a critical challenge for practical feasibility.
2.3.1 Individual fission (or thinning) for mixture models
In Gaussian mixture models, parameters are typically component-specific, meaning that they are assumed to be shared across all individuals within each component. As explained above and in Figure 1, this assumption poses a circular challenge as it requires knowledge of the true data structure to be able to accurately estimate the component-specific covariance matrices that are then needed to perform data fission or data thinning. To address this limitation, we propose an alternative approach where each observation is modeled as a realization of its own Gaussian distribution, i.e. $p_{i}(\boldsymbol{x}_{i})=f\left(\boldsymbol{x_{i}}|\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}\right)$. Thus, observations are no longer assumed to be identically distributed. Consequently, the covariance matrix $\boldsymbol{\Sigma}_{i}$ is no longer specific to the components but instead to the individual observations. Despite this individual-level definition, two individuals drawn from the same (unknown) component are expected to have very close variance parameters. By bypassing the components in the definition of $\boldsymbol{\Sigma}_{i}$, this modeling strategy theoretically encompasses both the global null ($K=1$) and the mixture ($K\geq 1$) settings, and opens up individual-level data fission and data thinning. As highlighted above, variance estimation remains crucial for the practical feasibility of these methods. This new modeling assumes individual variances that still need to be known (ideally), or precisely estimated in real-life settings.
2.3.2 Non-parametric local variance estimation
[…]
$\widehat{\sigma^{2}}_{i}=\frac{\sum\limits_{j=1}^{n}w_{ij}\left(x_{i}-\widehat{m}_{i}\right)^{2}}{\left(\sum\limits_{j=1}^{n}w_{ij}\right)-1}$ (2)
where $w_{ij}$ are individual-weights and $\widehat{m}_{i}=\sum\limits_{j=1}w_{ij}x_{i}\bigg{/}\sum\limits_{j=1}^{n}w_{ij}$ are individual-specific weighted means. Ideally, $w_{ij}$ should be zero (or very small) for all the observations $x_{j}$ that are not in the same component as $x_{i}$.
The catch of the variance computation in (2) lies in determining each individual weight $w_{ij}$. Since it is essential for $w_{ij}$ to appropriately capture the proximity between observations, we opted for a kernel-based definition: $w_{ij}=K\left(x_{i}-x_{j}\right)$, where $K(u)=\frac{1}{\sqrt{2\pi}h}e^{-\frac{u^{2}}{2h^{2}}}$ represents the Gaussian kernel providing the proximity measurement between $x_{i}$ and $x_{j}$. This ensures that individual weights reflect the local relationships within the data. This kernel choice focuses on nearby points, emphasizing their influence on the weighted variance estimate. The parameter $h$ in the definition of $K$ serves as the bandwidth parameter, controlling the width of the kernel and, consequently, the neighborhood around each observation $x_{i}$ that contributes to its weighted variance estimate. A smaller $h$ results in more localized estimations, emphasizing nearby points, but can underestimate the variances. Conversely, a larger $h$ includes a broader range of observations, and an excessively large value might consider almost all observations in the variance estimation, yielding an estimator close to the full sample one $\mathbf{\widehat{\Sigma}}$. Therefore, an optimal choice for $h$ would involve considering only the observations in the same components of the underlying mixture. However, achieving this ideal scenario is impractical as it is equivalent to knowing the true components of the mixture.
Bandwidth calibration is a crucial step in any kernel method (Heidenreich et al.,, 2013). We propose to use individual-specific bandwidth $h_{i}=h(x_{i})$ to reduce the bias in kernel density estimation. To accomplish this, we first estimate the changepoint, $i^{*}$ in the spread of distances from observation $x_{i}$ to all other observations. This changepoint delineates the change in components of the mixture: observations with distances preceding this changepoint are deemed part of the same component as $x_{i}$, whereas those following it are considered part of different components. Subsequently, we define $h_{i}=|x_{i}-x_{i^{*}}|$ as the distance between $x_{i}$ and the observation $x_{i^{*}}$ that is furthest from $x_{i}$ before the break in the mixture. Thus, $h_{i}$ is determined in such a way that the individual bandwidth accommodates only the observations in the same component as $x_{i}$, ensuring its size is sufficient to encompass all observations of the component (Chacón and Duong,, 2020).
3 Results
3.1 Quantification of Type I error rate as a function of the bias in the variance estimation
Independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ is guaranteed solely when the true variance is used for decomposition. Substituting the true variance $\boldsymbol{\Sigma}_{i}$ with an estimate $\widehat{\boldsymbol{\Sigma}}_{i}$ introduces correlations between these new random variables as demonstrated by Proposition 1 for data fission (a proof of this proposition could be found in Section $2$ of the Supplementary Materials) and by Proposition $10$ of Neufeld et al., 2023a .
Proposition 1
Let $\boldsymbol{X}\sim\mathcal{N}\left(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p\times p}\right)$ be a Gaussian random variable. Suppose we apply data fission on $\boldsymbol{X}$ as described in Table 1 but using $\boldsymbol{Z}\sim\mathcal{N}\left(\boldsymbol{0},\widehat{\boldsymbol{\Sigma}}\right)$, where $\widehat{\boldsymbol{\Sigma}}$ is an estimate of the $\boldsymbol{\Sigma}$. Then, $\operatorname{Cov}\left(\boldsymbol{X}^{(1)},\boldsymbol{X}^{(2)}\right)=\boldsymbol{\Sigma}-\widehat{\boldsymbol{\Sigma}}$.
Proposition 1 indicates that if $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ are not independent, as with a biased estimator of $\boldsymbol{\Sigma}$ that induces significant covariance, a split of a single component into two estimated clusters in $\boldsymbol{X}^{(1)}$ is easily transferred to $\boldsymbol{X}^{(2)}$ and results in false positives during the inference step (even if the latter is carried out on $\boldsymbol{X}^{(2)}$).
@@ 3.1 Quantification of Type I error rate as a function of the bias in the variance estimation
To further describe the repercussions of variance estimation in the context of data fission, we derived an analytical expression for the Type I error rate of the $t$-test as a function of the bias in estimating this parameter. Let $X_{1},\dots,X_{n}$ be $n$ independent and identically distributed random variables such that, for all $i=1,\dots,n$, $X_{i}\overset{i.i.d}{\sim}\mathcal{N}\left(\mu,\sigma^{2}\right)$. Here, $n$ represents the sample size. Recall that the sample $X_{1},\dots,X_{n}$ is normally distributed: it contains no real clusters, and therefore, no actual difference in means exists between subgroups of observations that is not a consequence of the clustering. Let $Z_{1},\dots,Z_{n}\overset{i.i.d}{\sim}\mathcal{N}\left(0,b^{2}\right)$ and $\tau\in,]0,+\infty[$. Our aim is to perform data fission of each $X_{i}$, using the $X^{(1)}_{i}$ for $k-$means clustering (with $K=2$) and the $X^{(2)}_{i}$ for differential testing between the two inferred clusters. Here, $b^{2}$ represents any value used as a plug-in for the variance of the $X_{i}$, and in particular, $b^{2}$ may be an estimate obtained for $\sigma^{2}$. Given the generation process of the $X_{i}$, it is established that regardless of the clustering on the $X^{(1)}_{i}$, there should be no mean difference between the estimated clusters on $X^{(2)}_{i}$ as long as independence is achieved. Let $C_{1}$ and $C_{2}$ be the two estimated clusters on the $X^{(1)}_{i}$ with the same intra-cluster variance, which is a reasonable hypothesis with $k-$means clustering as explained in Section $3$ of the Supplementary Materials. Since we are under the null hypothesis of no mean difference between the clusters, the $T$ statistic for the $t$-test between $C_{1}$ and $C_{2}$ using the $X^{(2)}_{i}$ is given by:
@@
where $\rho=\operatorname{Cor}\left(X^{(1)}_{i},X^{(2)}_{i}\right)=\frac{\left(\sigma^{2}-b^{2}\right)}{\sqrt{\left(\sigma^{2}+\tau^{2}b^{2}\right)\times\left(\sigma^{2}+\frac{1}{\tau^{2}}b^{2}\right)}}$. Section $3$ of the Supplementary Materials gives details on the derivation of this test statistic and its distribution. The associated Type I error rate for this test is given by $1-F(q_{\alpha/2})+F(-q_{\alpha/2})$, where $F$ is the cumulative distribution function of $\mathcal{N}\left(\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}},1\right)$ and $q_{\alpha/2}$ is the quantile of a standard Gaussian distribution $\mathcal{N}(0,1)$.
@@
We validated this theoretical result through numerical simulations, and conducted a detailed exploration of the influence of variance values and sample size on the resulting Type I error rate. We generated $n$ realizations of a Gaussian random variable with a mean $\mu=0$ and variances $\sigma^{2}$, with $1,\!000$ Monte Carlo repetitions. Subsequently, we used data fission with varying values of $b^{2}=\widehat{\sigma^{2}}$, obtaining $X^{(1)}$ for $k-$means clustering with $K=2$ and $X^{(2)}$ for testing mean differences between the two estimated clusters. Initially, we examine the impact of the original true variance $\sigma^{2}$ by considering $\sigma^{2}$ values of $\{0.01,0.25,1,4\}$ for a fixed sample size of $n=100$. Figure 3A, illustrates the relationship between the bias in estimating $\sigma^{2}$ and the Type I error rate, demonstrating a consistent agreement with the derived theoretical error rate.
We further documented the behavior of the Type I error for a fixed $\sigma^{2}=1$ with varying sample sizes $n\in\{50,100,200,100,500,1,\!000\}$ in Figure 3B, which shows the expected impact of the sample size on the Type I error rate. These findings collectively underscore the critical importance of accurately estimating the variance to achieve a well-calibrated Type I error rate with data fission.
@@
Figure 3: Impact of variance estimation on Type I error rate in data fission. Panel A: Evolution of the estimated Type I error rate in data fission as a function of the relative bias and the original variance of the data. Panel B: Evolution of the estimated Type I error rate in data fission as a function of the relative bias and the sample size.
3.2 Performances of the local variance estimator
We conducted simulation studies to assess the performance of the non-parametric variance estimator defined in (2). Generating univariate data akin to the motivating example in Figure 2 with $n=100$ realizations from a two-component univariate Gaussian mixture: $0.5\mathcal{N}\left(0,\sigma^{2}\right)+0.5\mathcal{N}\left(\delta,\sigma^{2}\right)$. We explored a range of ratio $\delta/\sigma$ values from 00 (i.e. no separation between the components, representing the global null of no clusters in the data) to $100$ (indicating an extreme separation). Different values of $\sigma^{2}$ were considered: $\sigma^{2}\in\{0.01,1,4\}$. For each pair $\left(\delta,\sigma^{2}\right)$ defining the mixture, variance was estimated from the data using our proposed weighted local variance, and the result was used for individual data fission. We then considered only the observations coming from the first component of the mixture for $k-$means clustering on the corresponding $X^{(1)}$ with $K=2$. This resulted in one true component splitting into two incorrect clusters. We then tested the mean differences between those two clusters using the $t$-test on $X^{(2)}$. This scenario was replicated $1,\!000$ times, and we computed the empirical Type I error rate at the $\alpha=5\%$ level.
Figure 4: Performance evaluation of the non-parametric variance estimator in a simulated univariate setting. All empirical results were obtained through $1,\!000$ simulations of the data. Panel A: Median relative bias, defined as $\left(\widehat{\sigma^{2}}-\sigma^{2}\right)/\sigma^{2}$ and its associated inter-quantile range, against the signal-to-noise ratio $\delta/\sigma$. Panel B: Type I error rate at the $\alpha=5\%$ level against $\delta/\sigma$. Panel C: Median relative bias and its associated inter-quantile range as a function of the sample size for three degrees of separation between the two components (informed by the ratio $\delta/\sigma$). Panel D: Type I error rate at the $\alpha=5\%$ level as a function of sample size.
Figure 4A shows that local variances are underestimated (manifested as a negative relative bias) until the signal vs. noise ratio $\delta/\sigma$ reaches approximately $3$ – a threshold value for separation in a Gaussian mixture model previously reported in the literature (Siffer et al.,, 2018; Hivert et al., 2024a, ). Consequently, there are no clear change-points in the spread of distances for individual observations and all observations should contribute significantly to the variance calculation. However, this is not the case here (with a bandwidth $h_{i}$ that is too small), leading to underestimated local variances. As the ratio increases within the range $3.5\leq\delta/\sigma\leq 10$, component separation becomes clearer but the identification of change-points remains challenging. Change-points detection is intricate for observations in the tails in-between the components, leading to non-zero weight for observations in both components. Consequently, an overestimation of the local variances (positive relative biases) is observed together with an increase in their associated inter-quantile range. Finally, for $\delta/\sigma>10$, sufficient component separation ensures consistent local variance estimation, highlighting the critical importance of clear-cut component separation for accurate variance estimation. So, as long as observations are well separated, our methodology outlined in 2.3.2 is able to provide accurate variance estimates. In Figure 4B, the Type I error rate at $5\%$ remains well calibrated for those values of $\delta/\sigma$ that ensure robust variance estimation (i.e. $\delta/\sigma>10$), illustrating the reliable testing performance with good component separations. Figure 4C demonstrates how increasing the sample size does not impact the performance of the non-parametric local variance estimations for three representative values of the ratio $\delta/\sigma\in\{0.5,5,10\}$, pointing towards component separation as the main driver of testing performance. The corresponding Type I error rates depicted in Figure 4D thus align with the observed relative bias results, underscoring the critical role of accurate estimation of local variances to ensure Type I error control.

### v3

@@ 3 Results
3.1 Limits in practical application of data fission and data thinning for mixture model
In the context of post-clustering inference on mixture models, data fission must be handled with care. The most straightforward approach is marginal data fission, which does not require any knowledge of the latent class memberships or mixture parameters. It relies solely on an estimate of the marginal covariance matrix, $\boldsymbol{\Sigma}$, typically approximated using empirical estimates. However, this simplicity comes at a cost: the two resulting random variables, $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$, are marginally uncorrelated as illustrated in Supplementary Table 1. This implies that the clustering structure present in the original data can only be preserved in one of the two new variables – never both – leading to a partial loss of information across the decomposition. Otherwise, if they both contain information of the latent clustering structure, they would be correlated.
Figure 1: Performance of marginal and conditional data fission in ideal and adverse clustering scenario over $1,000$ simulations under the Gaussian mixture setting. Panel A: Illustration of an ideal configuration where all parameters ($G$, $\pi_{g}$, $\boldsymbol{\mu}_{g}$, $\boldsymbol{\Sigma}_{g}$) are known. Panel B: Empirical statistical power (at the $5\%$ level) of the two-sample $t$-test between Cluster 11 and Cluster 22, as a function of $\tau$ and sample size. Panel C: Adjusted Rand Index (ARI) between $k$-means clustering and true labels after marginal or conditional data fission. Panel D: Illustration of an adverse scenario where a true component is erroneously split into two clusters. Panel E: Empirical Type I error rate (at the $5\%$ level) of the two-sample t-test under spurious Clusters 1 and 2, as a function of $\tau$ and sample size.
This limitation has practical consequences. Even in an ideal scenario where the true mixture parameters are known (Figure 1A), marginal data fission leads to poor transfer of clustering information from $\boldsymbol{X}_{i}^{(1)}$ to $\boldsymbol{X}_{i}^{(2)}$. As a result, downstream inference based on $\boldsymbol{X}_{i}^{(2)}$ suffers from a substantial loss in statistical power (Figure 1B), despite the clustering on $\boldsymbol{X}_{i}^{(1)}$ being nearly perfect (Figure 1C). The inability of $\boldsymbol{X}_{i}^{(2)}$ to benefit from the cluster structure learned in $\boldsymbol{X}_{i}^{(1)}$ stems directly from the marginal uncorrelation between the two fissioned random variables.
In more challenging scenarios where mixture parameters are unknown, marginal fission can even lead to misleading results. Figure 1D illustrates a case in which a single mixture component is wrongly split into two clusters due to imperfect clustering. Under marginal fission, the conditional covariance from equation (2) propagate this incorrect partition – and the artificial differences it introduces – into $\boldsymbol{X}_{i}^{(2)}$. As a result, false positives may arise when testing for differences between clusters that are, in truth, identical. This breakdown of Type I error control, shown in Figure 1E, stems from marginal fission’s conditional dependency property, allowing artifacts in $\boldsymbol{X}_{i}^{(1)}$ to contaminate inference in $\boldsymbol{X}_{i}^{(2)}$.
Conditional data fission offers a more principled alternative. Rather than ignoring the latent structure, it constructs two new random variables that are conditionally independent given the mixture component, while preserving the marginal between-cluster differences between the fissioned random variables. This ensures that clustering performed on $\boldsymbol{X}_{i}^{(1)}$ remains valid for inference in $\boldsymbol{X}_{i}^{(2)}$, without inducing spurious within-cluster correlations. In simulations, conditional fission consistently maintains higher clustering performances and statistical power (Figure 1B) and robust Type I error control (Figure 1E) when clustering is imperfect.
Nevertheless, conditional data fission is not without drawbacks. Its implementation requires knowledge of the component-specific covariance matrices and latent class memberships – quantities that are typically unknown in post-clustering contexts. In our simulations (Figure 1), conditional fission is applied using the true class memberships, providing a best-case benchmark. In real-world applications, however, this latent information must be first estimated through clustering, reintroducing the same post-clustering inference challenges that data fission aims to mitigate.
These simulations also highlighted the strong impact of $\tau$, which controls the proportion of information about $\boldsymbol{X}_{i}$ that is preserved in each fissioned datasets, for both the marginal and conditional data fission. When $\tau$ is small, most of the information is retained in $\boldsymbol{X}_{i}^{(1)}$, facilitating accurate clustering and yielding high Adjusted Rand Index values (Figure 1C). However, this comes at the cost of leaving little information in $\boldsymbol{X}_{i}^{(2)}$, especially under marginal fission, where clustering-related signal is not preserved. This mismatch leads to a sharp drop in statistical power when testing for differences between inferred clusters. Conversely, when $\tau$ is large, more signal is injected into $\boldsymbol{X}_{i}^{(2)}$, improving the potential for downstream testing, but weakening the clustering quality in $\boldsymbol{X}_{i}^{(1)}$. As a result, the clusters compared during testing may not reflect the true latent classes, again reducing power. Therefore, under marginal fission, power is maximized only at intermediate values of $\tau$, where the trade-off between clustering accuracy and test informativeness is partially balanced – but never fully resolved due to the structural disconnect between fissioned random variables. Benefits of conditional data fission are hence particularly pronounced in low-$\tau$ regimes, where signal in $\boldsymbol{X}^{(2)}$ is weak but the preservation of the clustering structure is strengthened by the marginal covariance.
3.2 Type I error relation with bias in the variance estimation
The conditional or marginal independence properties between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ reported in Supplementary Table 1 hold strictly when the true variance matrix–either $\boldsymbol{\Sigma}$ or the component-specific $\boldsymbol{\Sigma}_{g}$–is used for the decomposition. In practice, these quantities are typically unknown and must be estimated from data.
Proposition 1
@@ 3.2 Type I error relation with bias in the variance estimation
Proposition 1 (see also Section 3 of the Supplementary Materials for a proof) shows that replacing the true covariance with an estimate in the data fission process can induce dependencies between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$. A similar result is established by Proposition 10 in Neufeld, Dharamshi, Gao & Witten (2024) for data thinning. In particular, if the estimate $\widehat{\boldsymbol{\Sigma}}$ is biased, it can create spurious correlations between the fissioned datasets, enabling artificial clustering structures–such as the erroneous splitting of a single mixture component into multiple clusters in $\boldsymbol{X}^{(1)}$–to carry over into $\boldsymbol{X}^{(2)}$. As a consequence, Type I error may be inflated during inference, even though testing is conducted exclusively on $\boldsymbol{X}^{(2)}$.
@@
To further elucidate the consequences of variance misestimation in data fission, we derived an analytical expression for the Type I error rate of the $t$-test as a function of the estimation bias. This provides theoretical insight into how variance estimation errors propagate through the data fission process and impact downstream statistical inference. Let $X_{1},\dots,X_{n}$ be $n$ independent and identically distributed random variables such that, for all $i=1,\dots,n$, $X_{i}\overset{\text{iid}}{\sim}\mathcal{N}\left(\mu,\sigma^{2}\right)$. Here, $n$ represents the sample size. Recall that the sample $X_{1},\dots,X_{n}$ is normally distributed: it contains no real clusters, and therefore, no actual difference in means exists between subgroups of observations that is not a consequence of the clustering. Let $W_{1},\dots,W_{n}\overset{\text{iid}}{\sim}\mathcal{N}\left(0,b^{2}\right)$ and $\tau\in,]0,+\infty[$. Our aim is to perform data fission of each $X_{i}$, using the $X^{(1)}_{i}$ for $k$-means clustering (with $K=2$) and the $X^{(2)}_{i}$ for differential testing between the two inferred clusters. Here, $b^{2}$ represents any value used as a plug-in for the variance of the $X_{i}$, and in particular, $b^{2}$ may be an estimate obtained for $\sigma^{2}$. Given the generation process of the $X_{i}$, it is established that regardless of the clustering on the $X^{(1)}_{i}$, there should be no mean difference between the estimated clusters on $X^{(2)}_{i}$ as long as independence is achieved. Let $C_{1}$ and $C_{2}$ be the two estimated clusters on the $X^{(1)}_{i}$ with the same intra-cluster variance, which is a reasonable hypothesis with $k$-means clustering as explained in Section $4$ of the Supplementary Materials. Since we are under the null hypothesis of no mean difference between the clusters, the $T$ statistic for the $t$-test between $C_{1}$ and $C_{2}$ using the $X^{(2)}_{i}$ is given by:
@@
where $\rho=\operatorname{Cor}\left(X^{(1)}_{i},X^{(2)}_{i}\right)=\frac{\left(\sigma^{2}-b^{2}\right)}{\sqrt{\left(\sigma^{2}+\tau^{2}b^{2}\right)\times\left(\sigma^{2}+\frac{1}{\tau^{2}}b^{2}\right)}}$. Section $4$ of the Supplementary Materials gives details on the derivation of this test statistic and its distribution. The associated Type I error rate for this test is given by $1-F(q_{\alpha/2})+F(-q_{\alpha/2})$, where $F$ is the cumulative distribution function of $\mathcal{N}\left(\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}},1\right)$ and $q_{\alpha/2}$ is the quantile of a standard Gaussian distribution $\mathcal{N}(0,1)$.
@@
We validated this theoretical result through numerical simulations, and conducted a detailed exploration of the influence of variance values and sample size on the resulting Type I error rate. We generated $n$ realizations of a Gaussian random variable with a mean $\mu=0$ and variances $\sigma^{2}$, with 1,0001,\!000 Monte Carlo repetitions. Subsequently, we used data fission with varying values of $b^{2}=\widehat{\sigma^{2}}$, obtaining $X^{(1)}$ for $k$-means clustering with $K=2$ and $X^{(2)}$ for testing mean differences between the two estimated clusters. Initially, we examine the impact of the original true variance $\sigma^{2}$ by considering $\sigma^{2}$ values of $\{0.01,0.25,1,4\}$ for a fixed sample size of $n=100$. Figure 2A, illustrates the relationship between the bias in estimating $\sigma^{2}$ and the Type I error rate, demonstrating a consistent agreement with the derived theoretical error rate.We further documented the behavior of the Type I error for a fixed $\sigma^{2}=1$ with varying sample sizes $n\in\{50,100,200,100,500,1,\!000\}$ in Figure 2B, which shows the expected impact of the sample size on the Type I error rate. These results show the critical importance of accurately estimating the variance to achieve a well-calibrated Type I error rate with data fission.
@@
Figure 2: Impact of variance estimation on Type I error in data fission. Type I error as a function of the relative bias colored by (A) true data variance, or (B) sample size.
This issue with variance estimation fully accounts for the conditional correlations observed under marginal data fission. Given class membership, the Gaussian assumption is satisfied, and the correct covariance structure is component-specific. However, marginal fission relies on a global covariance estimate, $\boldsymbol{\Sigma}$, which averages over the component-specific covariances $\boldsymbol{\Sigma}_{g}$, $g=1,\dots,G$. This averaging introduces bias and leads to a misspecified covariance structure. As a result, dependencies between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ emerge at the component level, thereby compromising inference. By contrast, conditional data fission–when implemented with accurate, component-specific variance estimates–preserves the intended conditional independence structure and is therefore better suited for mixture settings.